OpenStudy (anonymous):

What are the values of theta in the interval 0

OpenStudy (anonymous):

\[0\le \theta \le 360\] \[\tan \theta -\sqrt{3} =0\]

OpenStudy (tkhunny):

You can't find even one? You might want to give a 30-60-90 right triangle a really close look.

OpenStudy (anonymous):

Well, I know there's 60, but that's it. Should I just know the value of 60º, 120º, 240º, and 300 by heart, or is there another way?

OpenStudy (tkhunny):

Good, but you should have mentioned that you already had one solution. 60º is in Quadrant I. Can you think of another quadrant where the tangent has the same sign? II? III?, IV?

OpenStudy (anonymous):

III

OpenStudy (anonymous):

180<270, so 240.

OpenStudy (tkhunny):

Reference Angle 60º Quadrant I and III Done. Good work.